OpenStudy (anonymous):

MEDAL + FAN FOR RIGHT ANSWER!!!! a water tank is a rectangular prism that is 66 meters wide 33 meters long, 66 meters high. A solid metal box is 11 meters long, 33 meters high, and 55 meters high is sitting inside the tank. the tank is filled with water. find the volume of water that is in the tank

OpenStudy (lsdxmdmaddicthc):

Okay so first you want to find the volume of the water tank You only said it's 33 by 66 so I'm going to assume it's 3 dimensional with the following attributes : 33*33*66 = 71874 meters cubed Now find the volume of the metal box 11*33*55 = 19965 meters cubed Now you can subtract the volume of the metal box from the volume of the container. 71874 - 19965 = 51909 meters cubed The volume of the water in the tank is 51909 meters cubed

OpenStudy (anonymous):

ok i changed the question srry

OpenStudy (anonymous):

@lsDxMdmaddicThc